OpenStudy (anonymous):

What is the multiplicative identity of the complex number –3 + 5i? a. 0 b. 1 c. –3 + 5i d. 3 – 5i

OpenStudy (justnick09):

b? i think

OpenStudy (anonymous):

thanks boo <3

OpenStudy (the_fizicx99):

That reasoning lol.

OpenStudy (zzr0ck3r):

lets call the multiplicative identity a then it must be that (–3 + 5i)a=a(–3 + 5i)=(–3 + 5i)

OpenStudy (anonymous):

so d?

OpenStudy (zzr0ck3r):

(3 – 5i)(3 – 5i)=9-30i-25=-16-30i

OpenStudy (zzr0ck3r):

what number when you multiply another number by it you get that other number back? 6x=6 what is x?

OpenStudy (anonymous):

1

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

what is the additive identity? a+? = a